OpenStudy (anonymous):

Find the vertex of the graph of the function. f(x) = 4x^2 + 24x + 32

OpenStudy (radar):

You're off to a good start, the equation is already in standard form.

OpenStudy (anonymous):

u kno vertex is a point of extreme value

OpenStudy (anonymous):

(4,5)

OpenStudy (radar):

You can sketch this by locating a few key points. To do this, completing the square is one method:\[4x ^{2}+24x+32\] divide thru by 4 getting:\[x ^{2}+6x+8\]\[x ^{2}+6x +9+8-9=(x+3)^{2}-1=0\]

OpenStudy (radar):

The vertex occurs at the point P(-3,-1), The axis is the line x=-3 Calculate some additional points and graph.

OpenStudy (radar):

I believe the factor 4 which was removed when completing the square has to be also considered.

OpenStudy (anonymous):

Hey @babydoll332! to find the vertex, given a quadratic equation ax^2 + bx + c...u can use the formula ! then u calculate  then, just plug that value into your original equation to give us the y coordinate of the vertex! 4(-3)^2 + 24(-3) + 32 = -4! (-3,-4) is the vertex!